es Esercise 07.13 Check My Work eBook used the Internet, at least occasionally (Pew website, October 23, 2012). The results questions in the Pew Internet & American Life Project esked adults if they OOne of the o showed that 434 out of 478 adults aged 18-29 answered Yes; 741 out of 833 adults aged 30-49 answered Yes; 1058 out of 1644 adults aged 50 and over answered Yes. Round your answers to four decimal places. a. Develop a point estimate of the proportion of adults aged 18-29 who use the Internet b. Develop a point estimate of the proportion of adults aged 30-49 who use the Internet C. Develop a point estimate of the proportion of adults aged 50 and over who use the Internet d. Comment on any relationship between age and internet use that seems apparent. Select your answer e.Suppose your target population of interest is that of all adults (18 years of age and over). Develop an estimate of the proportion of that population who use the Internet.
Answer #1

Solution:-

a) point estimate of the proportion of adults aged 18 – 29 who use the Internet = 454/478 = 0.9498

b) point estimate of the proportion of adults aged 30 – 49 who use the Internet = 741/833 = 0.8896

c) point estimate of the proportion of adults aged 50 and over who use the Internet = 1058/1644 = 0.6436

d) younger adults are more likely to use the internet

e) estimate of the total proportion of the population who use the Internet = (454+741+1058) / (478+833+1644)

estimate of the total proportion of the population who use the Internet = 2253/2955 = 0.7624